sam is 12 years younger than paulo . one third of sam's age is equal to one fourth of paulo's age how old is sam?
Answer by richwmiller(17219)

s=p-12
s/3=p/4
p=4/3s
s=p-12
s=4/3s-12
multiply by 3
3s=4s-36
-s=-36
s=36
p=48
check
36/3=48/4
12=12
ok